A clever engineer designs a "sprong" that obeys the force law Fx=−q(x−xe)3, where xe is the equilibrium position of the end of the sprong and q is the sprong constant. For simplicity, we'll let xe=0m. Then Fx=−qx3. Find an expression for the potential energy of a stretched or compressed sprong.
Answer #1

from the relation between force and potential energy

f = -du/dx

u = -fdx

u = qx^3 dx

u = q(x^4/4)

u = 1/4*qx^4